# Pseudo-real quantum fields

### v1

@@ 4 Generalisation to higher spins
In the preceding sections, we have focussed on the simplest, spin-zero representation. We now turn our attention to higher-spin representations. Of particular interest is the coupling to gauge fields, and we begin from a well-studied theory composed of two complex scalar fields with an anti-Hermitian mass mixing [15, 17, 16]. Doing so, will provide an illustration of how pseudo-reality allows us to gauge non-Hermitian quantum field theories and to couple them to gravity consistently. Moreover, this two complex scalar field theory falls under the second approach to constructing non-Hermitian quantum field theories, constructed ab-initio from non-Hermitian composite operators. We will see, however, that pseudo-reality also allows to interpret this theory as the analytic continuation of a Hermitian theory.
4.1 Two complex scalar field theory
@@
We start with the following free field theory (see, e.g., Refs. [15, 17, 16]):
$\mathcal{L}=\partial_{\mu}\tilde{\Phi}^{{\dagger}}\partial^{\mu}\Phi-\tilde{\Phi}^{{\dagger}}M^{2}\Phi\;,$ (4.1)
@@
where the field $\Phi\equiv(\phi_{1},\phi_{2})$ is composed of two complex scalar fields $\phi_{1}$ and $\phi_{2}$. As the original Lagrangian, composed of $\Phi$ and $\Phi^{\dagger}$ is $PT$-symmetric and parity-pseudo-Hermitian, the dual field $\tilde{\Phi}^{{\dagger}}\equiv(\tilde{\phi}_{1}^{{\dagger}},\tilde{\phi}_{2}^{{\dagger}})$ is given via the parity transformation $\hat{\mathcal{P}}$
$\hat{\tilde{\Phi}}^{{\dagger}}(x)=\hat{\mathcal{P}}^{-1}\hat{\Phi}^{{\dagger}}(x_{P})\hat{\mathcal{P}}\>\Pi\;,\qquad\Pi=\begin{pmatrix}+1&0\\ 0&-1\end{pmatrix}\;.$ (4.2)
@@
Here, $\Pi$ is the parity matrix, reflecting the scalar and pseudo-scalar nature of the fields $\phi_{1}$ and $\phi_{2}$ respectively. The mass matrix is skew-symmetric and $\Pi$-pseudo-Hermitian
$M^{2}=\begin{pmatrix}m_{1}^{2}&m_{5}^{2}\\ -m_{5}^{2}&m_{2}^{2}\end{pmatrix}\;,\qquad M^{2^{\dagger}}=\Pi M^{2}\Pi^{-1}\>,$ (4.3)
with $m_{1}^{2}$, $m_{2}^{2}$ and $m_{5}^{2}$ real and positive.
@@
$\hat{\phi}_{j}=\frac{1}{\sqrt{2}}\Big{(}\hat{\chi}_{j1}+i\hat{\chi}_{j2}\Big{)}\;,\qquad\text{and}\qquad\hat{\tilde{\phi}}^{{\dagger}}_{j}=\frac{1}{\sqrt{2}}\Big{(}\hat{\tilde{\chi}}^{{\dagger}}_{j1}-i\hat{\tilde{\chi}}^{{\dagger}}_{j2}\Big{)}\;.$ (4.4)
Here, we take the fields $\hat{\chi}_{jk}\in\mathbb{C}$, following an approach similar to Ref. [14] (cf. the alternative approach of Ref. [24], where the field is decomposed into two real components). If the field components $\hat{\chi}_{jk}$ and $\hat{\tilde{\chi}}_{jk}^{{\dagger}}$ are unrelated to each other, then we have eight degrees of freedom, rather than the four that we would expect for two complex scalar fields. We recover four degrees of freedom by imposing pseudo-reality on the field components $\hat{\chi}_{jk}$, namely that
$\hat{\tilde{\chi}}^{{\dagger}}_{jk}(x)=\hat{\chi}_{jk}(x)\;.$ (4.5)
@@
^$\hat{\mathcal{L}}=\frac{1}{2}\sum_{j,k}\partial_{\mu}\hat{\chi}_{jk}\partial^{\mu}\hat{\chi}_{jk}-\frac{1}{2}m_{1}^{2}\left(\hat{\chi}_{11}^{2}+\hat{\chi}_{12}^{2}\right)-\frac{1}{2}m_{2}^{2}\left(\hat{\chi}_{22}^{2}+\hat{\chi}_{21}^{2}\right)-im_{5}^{2}\left(\hat{\chi}_{11}\hat{\chi}_{22}-\hat{\chi}_{12}\hat{\chi}_{21}\right)\;,$ (4.6)
@@
$\Box\hat{\chi}_{11}+m_{1}^{2}\hat{\chi}_{11}+im_{5}^{2}\hat{\chi}_{22}=0\;,$ (4.7a)
$\Box\hat{\chi}_{12}+m_{1}^{2}\hat{\chi}_{12}-im_{5}^{2}\hat{\chi}_{21}=0\;,$ (4.7b)
$\Box\hat{\chi}_{21}+m_{2}^{2}\hat{\chi}_{21}-im_{5}^{2}\hat{\chi}_{12}=0\;,$ (4.7c)
$\Box\hat{\chi}_{22}+m_{2}^{2}\hat{\chi}_{22}+im_{5}^{2}\hat{\chi}_{11}=0\;.$ (4.7d)
This theory is analytically continued from a Hermitian theory by taking $m_{5}^{2}\to im_{5}^{2}$, and we see immediately, as for the $i\phi^{3}$ theory discussed in Sec. 3, that the fields $\hat{\chi}_{jk}$ cannot be real. We now need only to ensure the self-consistency of the pseudo-reality of the fields $\hat{\chi}_{jk}$.
@@
$\hat{\tilde{\chi}}^{{\dagger}}_{11}$ $=\hat{\mathcal{P}}^{-1}\hat{\chi}^{\dagger}_{11}\hat{\mathcal{P}}=\hat{\chi}_{11}\;,$ (4.8a)
$\hat{\tilde{\chi}}^{{\dagger}}_{12}$ $=\hat{\mathcal{P}}^{-1}\hat{\chi}^{\dagger}_{12}\hat{\mathcal{P}}=\hat{\chi}_{12}\;,$ (4.8b)
$\hat{\tilde{\chi}}^{{\dagger}}_{21}$ $=-\hat{\mathcal{P}}^{-1}\hat{\chi}^{\dagger}_{21}\hat{\mathcal{P}}=\hat{\chi}_{21}\;,$ (4.8c)
$\hat{\tilde{\chi}}^{{\dagger}}_{22}$ $=-\hat{\mathcal{P}}^{-1}\hat{\chi}^{\dagger}_{22}\hat{\mathcal{P}}=\hat{\chi}_{22}\;.$ (4.8d)
@@
4.2 Gauged model
A number of attempts have been made in the literature to gauge similar scalar models that incorporate two complex scalar fields [25, 26]. There are two possibilities: (i) couple the gauge field to a Hermitian but non-conserved current or (ii) couple the gauge field to a conserved but non-Hermitian current. Method (i) leads to an inconsistent Maxwell equation and breaks the gauge invariance of the Lagrangian. Method (ii) is in contradiction with the reality of the gauge field. We will show in this section that method (ii) is consistent so long as the gauge field is taken to be a pseudo-real vector field.
@@ 4.2 Gauged model
$\mathcal{L}=-\frac{1}{4}F_{\mu\nu}\tilde{F}^{{\dagger}\mu\nu}-\frac{1}{2\xi}(\partial_{\mu}A^{\mu})(\partial_{\mu}\tilde{A}^{{\dagger}\mu})+\tilde{D}_{\mu}^{{\dagger}}\tilde{\Phi}^{{\dagger}}D^{\mu}\Phi-\tilde{\Phi}^{{\dagger}}M^{2}\Phi\;,$ (4.9)
@@
$D_{\mu}=\partial_{\mu}-ieA_{\mu}\;,\qquad\tilde{D}_{\mu}^{{\dagger}}=\partial_{\mu}+ie\tilde{A}_{\mu}^{{\dagger}}\;,$ (4.10)
Consider the operator-valued conserved current
$\hat{j}^{\nu}=ie\left[\hat{\tilde{\Phi}}^{{\dagger}}\overset{\rightarrow}{D^{\nu}}\hat{\Phi}-\hat{\tilde{\Phi}}^{{\dagger}}\overset{\leftarrow}{\tilde{D}^{\nu{\dagger}}}\hat{\Phi}\right]\;.$ (4.11)
@@
$\phantom{=}+\frac{1}{2}e^{2}\left(\hat{A}^{\nu}+\hat{\tilde{A}}^{\nu{\dagger}}\right)\left[\hat{\chi}_{11}^{2}+\hat{\chi}_{12}^{2}+\hat{\chi}_{21}^{2}+\hat{\chi}_{22}^{2}\right]\;.$ (4.12)
@@
$\hat{\tilde{A}}^{\mu{\dagger}}(x)=\hat{\mathcal{P}}^{-1}\hat{A}^{\nu{\dagger}}(x_{P})\hat{\mathcal{P}}\Pi_{\nu}^{\phantom{\nu}\mu}=\hat{A}^{\mu}(x)\;,$ (4.13)
@@
The following Maxwell equations are now all internally consistent:
$\Box\hat{A}^{\mu}-\left(1-\frac{1}{\xi}\right)\partial^{\nu}\partial_{\mu}\hat{A}^{\mu}=\hat{j}^{\mu}\;,$ (4.14a)
$\Box\hat{\tilde{A}}^{\mu}-\left(1-\frac{1}{\xi}\right)\partial^{\nu}\partial_{\mu}\hat{\tilde{A}}^{\mu}=\hat{\tilde{j}}^{\mu}\;,$ (4.14b)
$\Box\hat{A}^{\mu{\dagger}}-\left(1-\frac{1}{\xi}\right)\partial^{\nu}\partial_{\mu}\hat{A}^{\mu{\dagger}}=\hat{j}^{\mu{\dagger}}\,,$ (4.14c)
and the pseudo-reality condition (4.13) implies that the photon field $A_{\mu}$, although complex, possesses only two propagating degrees of freedom.
We anticipate that this consistent pseudo-Hermitian formulation of a U(1) gauge theory will allow to revisit existing discrepancies in the treatments of the Higgs mechanism in similar non-Hermitian field theories (cf. Refs. [14, 25, 27]). This conjecture may be checked in future work.
4.3 Coupling to gravity
Having successfully gauged the two complex scalar model in the previous section, it is now clear how to couple non-Hermitian field theories to spin-2 fields. Consider, e.g., the free scalar sector described above. We can couple this theory minimally to Einstein gravity through the matter action
$S=\int{\rm d}^{4}x\sqrt{-g}\left[g^{\mu\nu}\partial_{\mu}\tilde{\Phi}^{{\dagger}}\partial_{\nu}\Phi-\tilde{\Phi}^{{\dagger}}M^{2}\Phi\right]\;,$ (4.15)
@@ 4.3 Coupling to gravity
$T^{\mu\nu}=\frac{1}{2}\partial^{(\mu}\tilde{\Phi}^{{\dagger}}\partial^{\nu)}\Phi-g^{\mu\nu}\mathcal{L}\;,$ (4.16)
@@
$\tilde{R}^{{\dagger}}_{\mu\nu}-\frac{1}{2}\tilde{g}_{\mu\nu}^{{\dagger}}\tilde{R}^{{\dagger}}=R_{\mu\nu}-\frac{1}{2}g_{\mu\nu}R\;,$ (4.18)
@@
Most notably, the pseudo-reality of the curvature tensors (4.18) and the metric (4.19) is determined by the pseudo-Hermiticity properties of all non-Hermitian sectors to which they couple, as was the case for the vector field in the previous section. We leave further study of these pseudo-real geometries to future work, but we anticipate that the observations in this work pave the way to consistent formulations of non-Hermitian cosmological models.
5 Conclusions
We have presented a general framework within which Hermiticity puzzles in pseudo-Hermitian quantum field theories can be resolved. We refer to this property as pseudo-reality. We first introduced the concept of pseudo-reality for complex numbers and then generalized it to quantum fields, promoting a real field defined on the real axis to a pseudo-real field defined on a contour in the complex plane. This approach resolves a long-standing inconsistency in the minimal coupling of pseudo-Hermitian sectors to gauge fields. We have focused here on a U(1) gauge theory and indicated how pseudo-reality enables a consistent coupling to gravity. Interestingly, the pseudo-reality conditions depend on the map $\eta$ that determines the pseudo-Hermiticity of the full Hamiltonian. Thus, the pseudo-reality condition of a given bosonic field is determined by the non-Hermitian sectors to which it couples, making clear the subtleties of treating interacting pseudo-Hermitian quantum field theories. We leave in-depth study of such theories and their couplings to gravity to future work.

### v2

@@ IV Generalisation to higher spins
In the preceding sections, we have focussed on the simplest, spin-zero representation. We now turn our attention to higher-spin representations. Of particular interest is the coupling to gauge fields, and we begin from a well-studied theory composed of two complex scalar fields with an anti-Hermitian mass mixing Alexandre:2017foi ; Alexandre:2020gah ; Sablevice:2023odu . Doing so will provide an illustration of how pseudo-reality allows us to gauge non-Hermitian quantum field theories and to couple them to gravity consistently. Moreover, this two complex scalar field theory falls under the second approach to constructing non-Hermitian quantum field theories, constructed ab-initio from non-Hermitian composite operators. We will see, however, that pseudo-reality also allows to interpret this theory as the analytic continuation of a Hermitian theory.
IV.1 Two complex scalar field theory
@@
We start with the following free field theory (see, e.g., Refs. Alexandre:2017foi ; Alexandre:2020gah ; Sablevice:2023odu ):
$\mathcal{L}=\partial_{\mu}\tilde{\Phi}^{{\dagger}}\partial^{\mu}\Phi-\tilde{\Phi}^{{\dagger}}M^{2}\Phi\;,$ (71)
@@
where the field $\Phi\equiv(\phi_{1},\phi_{2})$ is composed of two complex scalar fields $\phi_{1}$ and $\phi_{2}$. As the original Lagrangian, composed of $\Phi$ and $\Phi^{\dagger}$ is $PT$ symmetric and parity-pseudo-Hermitian, the dual field $\tilde{\Phi}^{{\dagger}}\equiv(\tilde{\phi}_{1}^{{\dagger}},\tilde{\phi}_{2}^{{\dagger}})$ can be given via the parity transformation $\hat{\mathcal{P}}$
$\hat{\tilde{\Phi}}^{{\dagger}}(x)=\hat{\mathcal{P}}^{-1}\hat{\Phi}^{{\dagger}}(x_{P})\hat{\mathcal{P}}\>\Pi\;,\qquad\Pi=\begin{pmatrix}+1&0\\ 0&-1\end{pmatrix}\;$. (72)
@@
Here, $\Pi$ is the parity matrix, reflecting the scalar and pseudo-scalar nature of the fields $\phi_{1}$ and $\phi_{2}$ respectively. The mass matrix has skew-symmetric off-diagonals and is $\Pi$-pseudo-Hermitian, with
$M^{2}=\begin{pmatrix}m_{1}^{2}&m_{5}^{2}\\ -m_{5}^{2}&m_{2}^{2}\end{pmatrix}\;,\qquad M^{2^{\dagger}}=\Pi M^{2}\Pi^{-1}\>,$ (73)
where $m_{1}^{2}$, $m_{2}^{2}$ and $m_{5}^{2}$ real and positive.
@@
$\hat{\phi}_{j}=\frac{1}{\sqrt{2}}\Big(\hat{\chi}_{j1}+i\hat{\chi}_{j2}\Big)\;,\qquad\text{and}\qquad\hat{\tilde{\phi}}^{{\dagger}}_{j}=\frac{1}{\sqrt{2}}\Big(\hat{\tilde{\chi}}^{{\dagger}}_{j1}-i\hat{\tilde{\chi}}^{{\dagger}}_{j2}\Big)\;.$ (74)
Here, we take the fields $\hat{\chi}_{jk}\in\mathbb{C}$, following an approach similar to Ref. Mannheim:2018dur (cf. the alternative approach of Ref. Fring:2019hue , where the field is decomposed into two real components). If the field components $\hat{\chi}_{jk}$ and $\hat{\tilde{\chi}}_{jk}^{{\dagger}}$ are unrelated to each other, then we have eight degrees of freedom, rather than the four that we would expect for two complex scalar fields. We recover four degrees of freedom by imposing pseudo-reality on the field components $\hat{\chi}_{jk}$, namely that
$\hat{\tilde{\chi}}^{{\dagger}}_{jk}(x)=\hat{\chi}_{jk}(x)\;.$ (75)
@@
$\hat{\mathcal{L}}=\frac{1}{2}\sum_{j,k}\partial_{\mu}\hat{\chi}_{jk}\partial^{\mu}\hat{\chi}_{jk}-\frac{1}{2}m_{1}^{2}\left(\hat{\chi}_{11}^{2}+\hat{\chi}_{12}^{2}\right)-\frac{1}{2}m_{2}^{2}\left(\hat{\chi}_{22}^{2}+\hat{\chi}_{21}^{2}\right)-im_{5}^{2}\left(\hat{\chi}_{11}\hat{\chi}_{22}-\hat{\chi}_{12}\hat{\chi}_{21}\right)\;,$ (76)
@@
$\Box\hat{\chi}_{11}+m_{1}^{2}\hat{\chi}_{11}+im_{5}^{2}\hat{\chi}_{22}=0\;,$ (77a)
$\Box\hat{\chi}_{12}+m_{1}^{2}\hat{\chi}_{12}-im_{5}^{2}\hat{\chi}_{21}=0\;,$ (77b)
$\Box\hat{\chi}_{21}+m_{2}^{2}\hat{\chi}_{21}-im_{5}^{2}\hat{\chi}_{12}=0\;,$ (77c)
$\Box\hat{\chi}_{22}+m_{2}^{2}\hat{\chi}_{22}+im_{5}^{2}\hat{\chi}_{11}=0\;.$ (77d)
This theory is analytically continued from a Hermitian theory by taking $m_{5}^{2}\to im_{5}^{2}$, and we see immediately, as for the $i\phi^{3}$ theory discussed in Sec. III, that the fields $\hat{\chi}_{jk}$ cannot be real. We now need only to ensure the self-consistency of the pseudo-reality of the fields $\hat{\chi}_{jk}$.
@@
$\hat{\tilde{\chi}}^{{\dagger}}_{11}$ $=\hat{\mathcal{P}}^{-1}\hat{\chi}^{\dagger}_{11}\hat{\mathcal{P}}=\hat{\chi}_{11}\;,$ (78a)
$\hat{\tilde{\chi}}^{{\dagger}}_{12}$ $=\hat{\mathcal{P}}^{-1}\hat{\chi}^{\dagger}_{12}\hat{\mathcal{P}}=\hat{\chi}_{12}\;,$ (78b)
$\hat{\tilde{\chi}}^{{\dagger}}_{21}$ $=-\hat{\mathcal{P}}^{-1}\hat{\chi}^{\dagger}_{21}\hat{\mathcal{P}}=\hat{\chi}_{21}\;,$ (78c)
$\hat{\tilde{\chi}}^{{\dagger}}_{22}$ $=-\hat{\mathcal{P}}^{-1}\hat{\chi}^{\dagger}_{22}\hat{\mathcal{P}}=\hat{\chi}_{22}\;.$ (78d)
@@
IV.2 Gauged model
A number of attempts have been made in the literature to gauge similar scalar models that incorporate two complex scalar fields Alexandre:2018xyy ; Millington:2019dfn . There are two possibilities: (i) couple the gauge field to a Hermitian but non-conserved current or (ii) couple the gauge field to a conserved but non-Hermitian current. Method (i) leads to an inconsistent Maxwell equation and breaks the gauge invariance of the Lagrangian. Method (ii) is in contradiction with the reality of the gauge field. We will show in this section that method (ii) is consistent so long as the gauge field is taken to be a pseudo-real vector field.
@@ IV.2 Gauged model
$\mathcal{L}=-\frac{1}{4}F_{\mu\nu}\tilde{F}^{{\dagger}\mu\nu}-\frac{1}{2\xi}(\partial_{\mu}A^{\mu})(\partial_{\mu}\tilde{A}^{{\dagger}\mu})+\tilde{D}_{\mu}^{{\dagger}}\tilde{\Phi}^{{\dagger}}D^{\mu}\Phi-\tilde{\Phi}^{{\dagger}}M^{2}\Phi\;,$ (79)
@@
$D_{\mu}=\partial_{\mu}-ieA_{\mu}\;,\qquad\tilde{D}_{\mu}^{{\dagger}}=\partial_{\mu}+ie\tilde{A}_{\mu}^{{\dagger}}\;,$ (80)
Let us now consider the operator-valued conserved current
$\hat{j}^{\nu}=ie\left[\hat{\tilde{\Phi}}^{{\dagger}}\overset{\rightarrow}{D^{\nu}}\hat{\Phi}-\hat{\tilde{\Phi}}^{{\dagger}}\overset{\leftarrow}{\tilde{D}^{\nu{\dagger}}}\hat{\Phi}\right$]\;. (81)
@@
$\phantom{=}+\frac{1}{2}e^{2}\left(\hat{A}^{\nu}+\hat{\tilde{A}}^{\nu{\dagger}}\right)\left[\hat{\chi}_{11}^{2}+\hat{\chi}_{12}^{2}+\hat{\chi}_{21}^{2}+\hat{\chi}_{22}^{2}\right]\;.$ (82)
@@
$\hat{\tilde{A}}^{\mu{\dagger}}(x)=\hat{\mathcal{P}}^{-1}\hat{A}^{\nu{\dagger}}(x_{P})\hat{\mathcal{P}}\Pi_{\nu}^{\phantom{\nu}\mu}=\hat{A}^{\mu}(x)\;,$ (83)
@@
The following operator-valued Maxwell equations are now all internally consistent:
$\Box\hat{A}^{\nu}-\left(1-\frac{1}{\xi}\right)\partial^{\nu}\partial_{\mu}\hat{A}^{\mu}=\hat{j}^{\nu}\;,$ (84a)
$\Box\hat{\tilde{A}}^{\nu}-\left(1-\frac{1}{\xi}\right)\partial^{\nu}\partial_{\mu}\hat{\tilde{A}}^{\mu}=\hat{\tilde{j}}^{\nu}\;,$ (84b)
$\Box\hat{A}^{\nu{\dagger}}-\left(1-\frac{1}{\xi}\right)\partial^{\nu}\partial_{\mu}\hat{A}^{\mu{\dagger}}=\hat{j}^{\nu{\dagger}}\,$, (84c)
and the pseudo-reality condition (83) implies that the photon field $A_{\mu}$, although complex, possesses only two propagating degrees of freedom.
We anticipate that this consistent pseudo-Hermitian formulation of a U(1) gauge theory will allow to revisit existing discrepancies in the treatments of the Higgs mechanism in similar non-Hermitian field theories (cf. Refs. Mannheim:2018dur ; Alexandre:2018xyy ; Alexandre:2019jdb ). This conjecture may be checked in future work.
IV.3 Coupling to gravity
Having successfully gauged the two complex scalar model in the previous section, we speculate on how to couple non-Hermitian field theories to spin-2 fields. Consider, e.g., the free scalar sector described above. We can couple this theory minimally to Einstein gravity through the matter action
$S=\int{\rm d}^{4}x\sqrt{-g}\left[g^{\mu\nu}\partial_{\mu}\tilde{\Phi}^{{\dagger}}\partial_{\nu}\Phi-\tilde{\Phi}^{{\dagger}}M^{2}\Phi\right]\;,$ (85)
@@ IV.3 Coupling to gravity
$T^{\mu\nu}=\frac{1}{2}\partial^{(\mu}\tilde{\Phi}^{{\dagger}}\partial^{\nu)}\Phi-g^{\mu\nu}\mathcal{L}\;,$ (86)
@@
$\tilde{R}^{{\dagger}}_{\mu\nu}-\frac{1}{2}\tilde{g}_{\mu\nu}^{{\dagger}}\tilde{R}^{{\dagger}}=R_{\mu\nu}-\frac{1}{2}g_{\mu\nu}R\;,$ (88)
@@
Most notably, this construction requires the metric to be complex. While the idea of complex metrics is not new (see, e.g., Ref. Witten:2021nzp and references therein), a comprehensive analysis of the implications of pseudo-reality for complex metrics is beyond the scope of this work. Interestingly, however, the pseudo-reality of the curvature tensors (88) and the metric (89) is determined by the pseudo-Hermiticity properties of all non-Hermitian sectors to which they couple, as was the case for the vector field in the previous section. We leave further study of these pseudo-real geometries to future work, but we anticipate that the observations in this work pave the way to consistent formulations of non-Hermitian cosmological models.
V Conclusions
We have presented a general framework within which Hermiticity puzzles in pseudo-Hermitian quantum field theories can be resolved. We refer to this property as pseudo-reality. We first introduced the concept of pseudo-reality for complex numbers and then generalized it to quantum fields, promoting a real field defined on the real axis to a pseudo-real field defined on a contour in the complex plane. This approach resolves a long-standing inconsistency in the minimal coupling of pseudo-Hermitian sectors to gauge fields. We have focused here on a U(1) gauge theory and indicated how pseudo-reality may enable a consistent coupling to gravity. Interestingly, the pseudo-reality conditions depend on the map $\eta$ that determines the pseudo-Hermiticity of the full Hamiltonian. Thus, the pseudo-reality condition of a given bosonic field is determined by the non-Hermitian sectors to which it couples, making clear the subtleties of treating interacting pseudo-Hermitian quantum field theories. We leave in-depth study of such theories and their couplings to gravity to future work.